# Interpretable Multi-Source Data Fusion Through Latent Variable Gaussian Process

### v1

@@ 4.1 Overview
The thermal age hardening phenomenon is driven by the composition of the FeCrAl alloy, temperature and duration of the thermal aging process. The hardness change of the alloy at different stages of thermal aging process is measured and used as an indication of the embrittlement. With the nuance of thermal aging phenomenon in FeCrAl alloys, material designers have the task of optimizing the alloy composition for maximum oxidation resistance and minimum embrittlement. The levers material designer have at their disposal is the composition of the FeCrAl alloys including dopants that can suppress the formation of $\alpha^{{}^{\prime}}$ phases.
Figure 5: Schemactics of the oxidation behavior of the FeCrAl alloy
@@ 4.2 Data Description
As with all material discovery efforts towards developing a generalizable multi-source data fusion model, the very first step is to gather data from research literature (papers, patents, etc.) and open-source databases. From an extensive search, seven well-cited and verified external sources of data were identified and collected. Additionally, there was also elaborate in-house testing conducted with GE Research to understand the embrittlement behavior of FeCrAl alloys. The sources of data are summarized through Table 3. The in-house generated samples are tagged as ’GE Data’. The rest of the colelcted samples are tagged by the name of the respective first authors of the paper. A total of 380 data points have been collected.
@@
Name of Data Source
Number of Data Samples
Source of Data
GE Data
$178$
Ravi, Sandipp Krishnan, et al. "Elucidating precipitation in FeCrAl alloys through explainable AI: A case study." Computational Materials Science 230 (2023): 112440. [7]
Kobayashi
$33$
Kobayashi, Satoru, and Takayuki Takasugi. "Mapping of 475 C embrittlement in ferritic Fe–Cr–Al alloys." Scripta Materialia 63.11 (2010): 1104-1107. [27]
Han
$55$
Han, Wentuo, et al. "Effect of Cr/Al contents on the 475ºC age-hardening in oxide dispersion strengthened ferritic steels." Nuclear Materials and Energy 9 (2016): 610-615. [32]
Kim
$4$
Kim, Hyunmyung, et al. "400 C aging embrittlement of FeCrAl alloys: Microstructure and fracture behavior." Materials Science and Engineering: A 743 (2019): 159-167. [37]
Ejenstam
$12$
Ejenstam, Jesper, et al. "Microstructural stability of Fe–Cr–Al alloys at 450–550 C." Journal of Nuclear Materials 457 (2015): 291-297. [38]
Yang
$64$
Yang, Z., et al. "Aluminum suppression of $\alpha^{{}^{\prime}}$ precipitate in model Fe–Cr–Al alloys during long-term aging at 475 C." Materials Science and Engineering: A 772 (2020): 138714. [39]
Capdevila
$16$
Capdevila, Carlos, et al. "Phase separation in PM 2000™ Fe-base ODS alloy: Experimental study at the atomic level." Materials Science and Engineering: A 490.1-2 (2008): 277-288. and Capdevila, Carlos, Michael K. Miller, and Jesús Chao. "Phase separation kinetics in a Fe–Cr–Al alloy." Acta materialia 60.12 (2012): 4673-4684. [29, 40]
Pinkas
$18$
Pinkas, Malki, et al. "Sensitivity of thermo-electric power measurements to $\alpha$$\alpha^{{}^{\prime}}$ phase separation in Cr-rich oxide dispersion strengthened steels." Journal of Materials Science 50 (2015): 4629-4635. [41]
Total
$380$
@@ 4.3 Model Development
To contrast and compare multi-source modeling capabilities, GP and LVGP models are trained on the collected data. The inputs for the GP model are the composition of the alloy constituents (Iron, Aluminium, Chromium, Molbidium, Zirconium, Manganese, Silicon, Tungsten, Yttrium and Titanium) defined as weight percentages, temperature and duration of the thermal aging process, totalling up to twelve (12) unique inputs. The output that is used to quantify the embrittlement behavior is represented through the measured hardness change (HV). On the other side, the inputs for the LVGP model contain an additional qualitative input that parameterizes each source of data as a distinct categorical level (option) along with all the other 12 numerical inputs used for the GP model. The output for the LVGP is identical to the one used for the GP model.
@@ 4.4 Results and Discussion
Within this study, (1) the generalizability of the models, and (2) the performances of the models with respect to the experimental in-house information source that can be controlled is investigated. The main goal of (1) is to observe the usefulness of the models when one has control to multiple sources of information to perform further experimentations, whereas the main goal of (2) is to observe which of the models is more reliable when one has further control over a source of information. Out of all 380 available samples, 344 samples is used to observe the generazability of models through cross validation (CV) and set aside 36 GE Data points as a seperate testing set to demonstrate the prediction capabilities of the models with respect to the information source that can be control and perform further experiments on. For the remaining 344 samples, 5-fold cross validation (around 275 training and 69 validation samples per fold) is employed to evaluate the generalizability of both models. For CV tasks in LVGP, stratified sampling is used to make sure datasets used for training and testing represent the overall distribution of the information sources. Then, both models are trained using the 344 samples and tested on the separate 36 GE testing data. The training, validation and testing samples were held identical for both models. The results of this study are shown in Section 4.4.1. In addition to comparing two ML models on the multi-source data, LVGP is further compared with a seperate GP model trained only on GE Data. For ease of notation,this GP model is denoted as GP-GE. GP-GE was only trained on GE Data samples available in the remaining 344 sample set (142 samples). Both LVGP and GP-GE are tested on the same sepearate 36 GE Data samples. The results of this study are shown in Section 4.4.2.
@@
Furtheremore, the LVGP model is investigated to extract physical insights and to obtain interpretability regarding the sources of information and the system of interest through the latent variables. The results of this examination is highlighted in Section 4.4.3. Finally, based on insights obtained from the latent space, a targeted source selection through data filtering mechanism is implemented to improve the predictive capabilities of multi-source data fusion in Section 4.4.4.
4.4.1 Multi-Source Data Fusion Comparision: LVGP and GP
@@
The predictive capability comparision between multi-source data fusion through GP and LVGP models are summarized in Table 4 (Model Types: GP and LVGP) and Panels A and B of Figure 7. From the summarized prediction metric, NRMSE (Table 4), we observe that the altough GP model perfoms slightly better in training, we observe a clear improvement in the cross validation (CV), which contains the samples that the model has not seen during training. The parity plots from the CV study are shown in Supplementary Figure 1. The CV results demonstrate that incorporating information sources as a seperate variable through LVGP can lead to more comprehensive and generalizable prediction model. Looking further at the testing set which contains seperate 36 GE Data samples, LVGP performs significantly better compared to the GP model, demonstrating that expliction categorization of the sources provides much accurate source-specfic predictions. Furthermore, the 45 degree parity plot between the GP and LVGP model across the training and testing data (first row of Figure 6) demonstrate that in addition to delivering accurate predictions, LVGP provides lower prediction uncertainities, which can be inferred from the error bars.
@@
Table 4: Predictive performance comparision for GP, LVGP, GP-GE, and LVGP-Target models on FeCrAl alloys
Model   Type
Mean Training NRMSE
Mean CV NRMSE
Testing NRMSE
GP
$0.060$
$0.105$
$0.088$
LVGP
$0.064$
$0.100$
$0.066$
GP-GE
$0.057$
LVGP-Target
$0.053$
@@
To strengthen our case of using multi-source modeling through LVGP, response surface plots from both the GP and LVGP models (with GE Data as reference) are also provided through Panels A and B of Figure 6. The second and third row depict the mean and standard deviation of the hardness changes across variation of the Cr (12%-24%) and Al (0%-7%) and the fourth row depicts the combined mean and standard deviation in a 3D representation. The model validity is assessed by focusing on the mean and variation behavior of the response surface. It is noted that the response surface of the hardness changes are highly undulated with sharp peaks and valleys with the GP model. On the other hand, using the GE source as ground truth for prediction, the response surface is much smoother and more well-behaved with the LVGP model. One can argue that the unsmooth behavior is caused by the varying level of noises coming from different sources in the data set when fitting all of them to one single GP model. On the other hand, LVGP has the capability of smoothing out the curves from multiple sources after the introduction of the latent variables. From a material science perspective, it is very unlikely that phenomenon driving the hardness changes across the composition are highly transient and acute. This important observation highlights the need of accounting for known and unknown underlying physcial parameters when data from numerous sources are combined for data fusion. The samples collected from different papers might have had differences in manufacturing process, thermal aging procedure, and testing technique that are not accounted for in a single source model. Furthermore, the downstream tasks of optimization, sensitivity analysis, uncertainty quantification would also be severely affected by such distinctions and sharp undulations of the surrogate models, thus making the case of employing multi-source modeling through LVGP once again.
4.4.2 Comparision of Multi-Source LVGP with Single-Source GP-GE
@@
A possible argument to achieve more accurate ML models is to build models only using the data source that one has control over, which would be the GE Data in the current study. To demonstrate the advantages of using multi-source data fusion through LVGP over a single-source model for the specified source of interest, the prediction performances of LVGP and the GP model trained on the GE Data only, which was previously denoted as GP-GE, are compared. Table 4 (Model Types: LVGP and GP-GE) and Panels B and C of Figure 6 demonstrate the prediction metrics and the response surface obtained from the models. Due to the differences in the data used to train models (multi-source vs single-source), testing accuracy of the models are studied on the common 36 GE Data testing set. The NMRSE prediction metrics from Table 4 demonstrate that incorporating other available sources into the modeling yield very similar predictive performances, with a slightly better results towards GP-GE model. However, the comparing the 45-degree parity plots between the two models (first row of Figure 6), we observe that the prediction uncertainities provided by the LVGP are much smaller compared to the prediction uncertainities provided by the GP-GE model. A possible explanation for this phenomenon could be that the knowledge incorporated from other sources contributes to the learning of the material system and therefore increases the confidence in the model when making predictions. Furthermore, the response surfaces between the two models show high similarity, indicating that the LVGP was able to capture the response surface of the specific source even when other data sources exist. The final important observation is that the predictions provided by the LVGP model possess larger and lower maximum values with lower uncertainities at different locations. This result could be significant and advantageous for materials design applications as it can give a multi-step headstart to the design optimization task through more confident and accurate predictions.
@@
Figure 6: Predictive performance comparision between (a) GP, (b) LVGP, and (c) GP-GE models. First row shows the parity plots of the models with respect to their respective training and GE Data testing sets. The x and y axis show the true and predicted hardness change (HV), respectively, and error bars demonstrate the prediction uncertainity associated with each prediction. Second, third and fourth row show the Mean, Standard Deviation and 3D Response Surfaces of the models at 12%<Cr<24%, 0%<Al<7%, 88% <Fe<69%, Mo = 0%, Zr = 0%, Mn = 0%, Si = 0%, W = 0%, Y = 0%, Ti = 0%, Temperature = ${}^{\circ}$ and Duration = 5000 hours, respectively
@@
For research transparency, it is important to note that without further experimental testing and validation in the highlighted compositional space, it cannot be stated which model is a true and generalizable model. The goal of the current work is to illustrate the existance of data discrepancies that could harm the multi-source modeling and how they can be integrated into modeling using LVGP. However, the current research study demonstrates that using LVGP as a multi-source data fusion model can be advantegeous due to providing (1) smoother response surfaces closer to real physical material behaviour, and (2) accurate and confident predictions which can be highly beneficial for downstream tasks such as material optimization, sensitivity analysis, and uncertainty quantification.
@@ 4.4.3 Interpretability From The Latent Variables
In addition to providing smoother response surfaces closer to the physical behaviour, the LVGP model provides interpretability into different sources through the latent space and the quantified dissimilarity metric, $D$. The obtained latent variables are shown in shown in Figure 7a. As previously highlighted, the spatial relations (distances) between the information sources can reveal certain information and provide further interpretability and insights on the sources. For analysis purposes, the GE Data source is used as reference source and compared with rest of the sources as GE Data source can be re-manufactured, optimized, validated, and tested further. It is observed from the latent space that the placement of the sources vary significantly with respect to the reference source. Specifically, the top three sources of data that are dissimilar (higher $D$ value) from the GE Data source are Yang, Kobayashi, and Pinkas, whereas the top three sources of data that are similar (lower $D$ values) to the GE Data source are Han, Kim, and Ejenstam.
@@
The spatial differences in the latent space can be attributed to differences in the input space, variations in the output space, or underlying fundamental differences in the input-output relationships that cannot be captured from the data. To investigate and interpret the latent spaces, a deeper study on the input and output (property) space is conducted. First, a principle component analysis (PCA) is performed to observe the differences in the input space. Figure 7c shows that the two principle components of the GE Data source cover and extend beyond the principle components of the other sources, indicating that the neither the closeness or the remoteness of the sources to the GE Data source in the latent space can be attributed to the differences in the input space. Looking further into the output distributions (Figure 7d), it is seen that the GE Data once again covers and extends beyond the properties of other available sources, demonstrating that the differences in the latent space is not due to the variabilities in the output space. As a result, the most impactful observation made from the latent space of data sources is that there are some fundamental differences between the data sources that are not explicitly known. As mentioned earlier, these possible differences for the current case can be due to differences in the manufacturing process of the alloy, testing methods, measuring procedure, underlying phenomenon, and more. Thus, it can be concluded that important differences exists between the information sources and LVGP is not only able to capture these variabilities but also learns the relationships between the sources.
@@
A further validation study is done by seperating the GE Data into two different sources as GE Data 1 and GE Data 2 and retraining the LVGP model to observed the latent space. The aim of this validation study is to observe the new latent space when the same underlying experimental procedures are implemented for two resources along with other sources that have different unknown underlying procedures. The new constructed latent space is shown in Figure 7b. A significant observation from the validation study is that the latent variables of GE Data 1 and GE Data 2 are located adjacent to each other. This result was expected since same experimental procedures are implemented for both of the resources. However, the results also once again demonstrated and validated that the LVGP model is able to capture the differences in the known and unknown underlying parameters that influence the material system and learn the relationships between the sources. This remarkable capability provided by the LVGP can be highly important for further downstream tasks including but not limited to model validation, dimension reduction, sensitivity analysis, and more. It is important to highlight that without using the LVGP approach proposed in this paper, it would be difficult to learn and assess the relationships between the data sources.
@@
Figure 7: (a) Latent Space of FeCrAl sources from LVGP. GE Data is denoted as the reference source (*) for dissimilarity metric, $D$, calculations. (b) Latent Space obtained from the validation study on interpretation. GE Data 1 is denoted as the reference source (*). (c) Two principal components of the input space and (d) output distribution of HV with respect to information sources
@@
4.4.4 Enchancing Modeling Capabilities Through Targeted Source Selection Based on Latent Variables

### v3

@@ 4.1 Overview
The thermal age hardening phenomenon is driven by the composition of the FeCrAl alloy, temperature, and duration of the thermal aging process. The hardness change of the alloy at different stages of thermal aging process is measured and used as an indication of the embrittlement. With the nuance of thermal aging phenomenon in FeCrAl alloys, material designers have the task of optimizing the alloy composition for maximum oxidation resistance and minimum embrittlement. The levers that material designers have at their disposal is the composition of the FeCrAl alloys including dopants that can suppress the formation of $\alpha^{{}^{\prime}}$ phases.
Figure 5: Schematic of the oxidation behavior of the FeCrAl alloy
@@ 4.2 Data Description
As with all material discovery efforts towards developing a generalizable multi-source data fusion model, the very first step is to gather data from research literature (papers, patents, etc.) and open-source databases. From an extensive search, seven well-cited and verified external sources of data were identified and collected. Additionally, there was also elaborate in-house testing conducted with GE Research to understand the embrittlement behavior of FeCrAl alloys. The sources of data are summarized through Table 3. The in-house generated samples are tagged as ’GE Data’. The rest of the collected samples are denoted by the name of the respective first authors of the paper. A total of 380 data points has been collected.
@@
Name of Data Source Number of Data Samples Source of Data
GE Data $178$ Ravi, Sandipp Krishnan, et al. "Elucidating precipitation in FeCrAl alloys through explainable AI: A case study." Computational Materials Science 230 (2023): 112440. [ravi2023elucidating]
Kobayashi $33$ Kobayashi, Satoru, and Takayuki Takasugi. "Mapping of 475 C embrittlement in ferritic Fe–Cr–Al alloys." Scripta Materialia 63.11 (2010): 1104-1107. [kobayashi2010mapping]
Han $55$ Han, Wentuo, et al. "Effect of Cr/Al contents on the 475ºC age-hardening in oxide dispersion strengthened ferritic steels." Nuclear Materials and Energy 9 (2016): 610-615. [han2016effect]
Kim $4$ Kim, Hyunmyung, et al. "400 C aging embrittlement of FeCrAl alloys: Microstructure and fracture behavior." Materials Science and Engineering: A 743 (2019): 159-167. [kim2019400]
Ejenstam $12$ Ejenstam, Jesper, et al. "Microstructural stability of Fe–Cr–Al alloys at 450–550 C." Journal of Nuclear Materials 457 (2015): 291-297. [ejenstam2015microstructural]
Yang $64$ Yang, Z., et al. "Aluminum suppression of $\alpha^{{}^{\prime}}$ precipitate in model Fe–Cr–Al alloys during long-term aging at 475 C." Materials Science and Engineering: A 772 (2020): 138714. [yang2020aluminum]
Capdevila $16$ Capdevila, Carlos, et al. "Phase separation in PM 2000™ Fe-base ODS alloy: Experimental study at the atomic level." Materials Science and Engineering: A 490.1-2 (2008): 277-288. and Capdevila, Carlos, Michael K. Miller, and Jesús Chao. "Phase separation kinetics in a Fe–Cr–Al alloy." Acta materialia 60.12 (2012): 4673-4684. [capdevila2008phase, capdevila2012phase]
Pinkas $18$ Pinkas, Malki, et al. "Sensitivity of thermo-electric power measurements to $\alpha$$\alpha^{{}^{\prime}}$ phase separation in Cr-rich oxide dispersion strengthened steels." Journal of Materials Science 50 (2015): 4629-4635. [pinkas2015sensitivity]
Total $380$
@@ 4.3 Model Development
To contrast and compare multi-source modeling capabilities, GP and LVGP models are trained on the collected data. The inputs for the GP model are the composition of the alloy constituents (Iron, Aluminum, Chromium, Molybdenum, Zirconium, Manganese, Silicon, Tungsten, Yttrium and Titanium) defined as weight percentages (wt%), temperature, and duration of the thermal aging process, totaling up to twelve (12) unique inputs. The output that is used to quantify the embrittlement behavior is represented through the measured hardness change (HV). On the other side, the inputs for the LVGP model contain an additional qualitative input that parameterizes each source of data as a distinct categorical level (option) along with all the other 12 numerical inputs used for the GP model. The output for the LVGP is identical to the one used for the GP model.
@@ 4.4 Results and Discussion
Within this study, (1) the generalizability of the models, and (2) the performances of the models with respect to the experimental in-house information source that can be controlled is investigated. The main goal of (1) is to observe the usefulness of the models when one has control to multiple sources of information to perform further experimentations, whereas the main goal of (2) is to observe which of the models is more reliable when one has further control over a source of information. Out of all 380 available samples, 344 samples are used to observe the generalizability of models through cross validation (CV) and set aside 36 GE Data points as a separate testing set to demonstrate the prediction capabilities of the models with respect to the information source that can be control and perform further experiments on. For the remaining 344 samples, 5-fold cross validation (around 275 training and 69 validation samples per fold) is employed to evaluate the generalizability of both models. For CV tasks in LVGP, stratified sampling is used to make sure datasets used for training and testing represent the overall distribution of the information sources. Then, both models are trained using the 344 samples and tested on the separate 36 GE testing data. The training, validation and testing samples were held identical for both models. The results of this study are shown in Section 4.4.1. In addition to comparing two ML models on the multi-source data, LVGP is further compared with a separate GP model trained only on GE Data. For ease of notation, this GP model is denoted as GP-GE. GP-GE was only trained on GE Data samples available in the remaining 344 sample set (142 samples). Both LVGP and GP-GE are tested on the same separate 36 GE Data samples. The results of this study are shown in Section 4.4.2.
@@
Furthermore, the LVGP model is investigated to extract physical insights and to obtain interpretability regarding the sources of information and the system of interest through the latent variables. The results from this examination are highlighted in Section 4.4.3. Finally, based on insights obtained from the latent space, a targeted source selection through data filtering mechanism is implemented to improve the predictive capabilities of multi-source data fusion in Section 4.4.4.
4.4.1 Multi-Source Data Fusion Comparison: LVGP and GP
@@
The predictive capability comparison between multi-source data fusion through GP and LVGP models are summarized in Table 4 (Model Types: GP and LVGP) and Panels A and B of Figure 8. From the summarized prediction metric, NRMSE (Table 4), although GP model performs slightly better in training instances, we observe a clear improvement in the cross validation (CV) metric using LVGP, which contains the samples that the model has not seen during training. The parity plots from the CV study are shown in Supplementary Figure 1. The CV results demonstrate that incorporating information sources as a separate variable through LVGP can lead to a more comprehensive and generalizable prediction model. Looking further at the testing set which contains separate 36 GE Data samples, LVGP performs significantly better compared to the GP model, demonstrating that explicit categorization of the sources provides more accurate source-specific predictions. Furthermore, the 45-degree parity plot between the GP and LVGP model across the training and testing data (first row of Figure 6) demonstrate that in addition to delivering accurate predictions, LVGP provides lower prediction uncertainties, which can be inferred from the error bars.
@@
Table 4: Predictive performance comparison between GP, LVGP, GP-GE, and LVGP-Target models on FeCrAl alloys
Model   Type Mean Training NRMSE Mean CV NRMSE Testing NRMSE
GP $0.060$ $0.105$ $0.088$
LVGP $0.064$ $0.100$ $0.066$
GP-GE $0.057$
LVGP-Target $0.053$
@@
To strengthen our case of using multi-source modeling through LVGP, response surface plots from both the GP and LVGP models (with GE Data as reference) are also provided through Panels A and B of Figure 6. The second and third row depict the mean and standard deviation of the hardness changes across variation of the Cr (12%-24%) and Al (0%-7%) and the fourth row depicts the combined mean and standard deviation in a 3D representation. The model validity is assessed by focusing on the mean and variation behavior of the response surface. It is noted that the response surface of the hardness changes are highly undulated with sharp peaks and valleys with the GP model. On the other hand, using the GE source as ground truth for prediction, the response surface is much smoother and more well-behaved with the LVGP model. One can argue that the unsmooth behavior is caused by the varying level of noises coming from different sources in the data set when fitting all of them to one single GP model. On the other hand, LVGP has the capability of smoothing out the curves from multiple sources after the introduction of the latent variables. From a material science perspective, it is very unlikely that phenomenon driving the hardness changes across the composition are highly transient and acute. This important observation highlights the need of accounting for known and unknown underlying physical parameters when data from numerous sources are combined for data fusion. The samples collected from different papers might have had differences in manufacturing process, thermal aging procedure, and testing technique that are not accounted for in a single source model. Furthermore, the downstream tasks of optimization, sensitivity analysis, uncertainty quantification would also be severely affected by such distinctions and sharp undulations of the surrogate models, thus making the case of employing multi-source modeling through LVGP once again. Finally, LVGP has the capability of providing source-specific predictions, which is not plausible with a single GP model that cannot differentiate between the sources. The response surface predictions provided by LVGP model for the remaining sources of information are provided in Supplementary Figure 7. This property provided by the LVGP can be highly advantageous for the cases when one has control over multiple sources of information.
4.4.2 Comparison of the Multi-Source LVGP with the Single-Source GP-GE
@@
A possible argument to achieve more accurate ML models is to build models only using the data source that one has control over, which would be the GE Data in the current study. To demonstrate the advantages of using multi-source data fusion through LVGP over a single-source model for the specified source of interest, the prediction performances of LVGP and the GP model trained on the GE Data only, which was previously denoted as GP-GE, are compared. Table 4 (Model Types: LVGP and GP-GE) and Panels B and C of Figure 6 demonstrate the prediction metrics and the response surface obtained from the models. Due to the differences in the data used to train models (multi-source vs single-source), testing accuracy of the models are studied on the common 36 GE Data testing set. The NMRSE prediction metrics from Table 4 demonstrate that incorporating other available sources into the modeling yield very similar predictive performances, with slightly better results towards GP-GE model. However, comparing the 45-degree parity plots between the two models (first row of Figure 6), we observe that the prediction uncertainties provided by the LVGP are much smaller compared to the prediction uncertainties provided by the GP-GE model. A possible explanation for this phenomenon could be that the knowledge incorporated from other sources contributes to the learning of the material system and therefore increases the confidence in the model when making predictions. Furthermore, the response surfaces between the two models show high similarity, indicating that the LVGP was able to capture the response surface of the specific source even when other data sources exist. The final very important observation is that the predictions provided by the LVGP model possess larger and lower maximum mean prediction values with lower uncertainties (standard deviation predictions) at different locations compared to both GP and GP-GE models. This observation demonstrates that the LVGP is able can perform extrapolation with much higher confidence by learning from other sources. This result could be significant and advantageous for materials design applications as it can give a multi-step head start to the design optimization task through more confident and accurate predictions.
@@
Figure 6: Predictive performance comparison between (a) GP, (b) LVGP, and (c) GP-GE models. The first row shows the parity plots of the models with respect to their respective training and GE Data testing sets. The x and y axis show the true and predicted hardness change (HV), respectively, and error bars demonstrate the prediction uncertainty associated with each prediction. Second, third and fourth row show the Mean, Standard Deviation and 3D Response Surfaces of the models at 12%<Cr<24%, 0%<Al<7%, 88% <Fe<69%, Mo = 0%, Zr = 0%, Mn = 0%, Si = 0%, W = 0%, Y = 0%, Ti = 0%, Temperature = 475C and Duration = 5000 hours, respectively
Figure 7: Source-specific 3D response surfaces provided by LVGP for the remaining FeCrAl alloy information sources at 12%<Cr<24%, 0%<Al<7%, 88% <Fe<69%, Mo = 0%, Zr = 0%, Mn = 0%, Si = 0%, W = 0%, Y = 0%, Ti = 0%, Temperature = 475C and Duration = 5000 hours, respectively. The colorbar represents the prediction uncertainity as standard deviation.
@@
For research transparency, it is important to note that without further experimental testing and validation in the highlighted compositional space, it cannot be stated which model is a true and generalizable model. The goal of the current work is to illustrate the existence of data discrepancies that could harm multi-source modeling and how they can be integrated into modeling using LVGP. However, the current research study demonstrates that using LVGP as a multi-source data fusion model can be advantageous due to providing (1) smoother response surfaces closer to real physical material behavior, and (2) accurate and confident predictions which can be highly beneficial for downstream tasks such as material optimization, sensitivity analysis, and uncertainty quantification.
@@ 4.4.3 Interpretability from the Latent Variables
In addition to providing smoother response surfaces closer to the physical behavior, the LVGP model provides interpretability into different sources through the latent space and the quantified dissimilarity metric, $D$. The obtained latent variables are shown in shown in Figure 8a. As previously highlighted, the spatial relations (distances) between the information sources can reveal certain information and provide further interpretability and insights on the sources. For analysis purposes, the GE Data source is used as reference source and compared with rest of the sources as GE Data source can be re-manufactured, optimized, validated, and tested further. It is observed from the latent space that the placement of the sources varies significantly with respect to the reference source. Specifically, the top three sources of data that are dissimilar (higher $D$ value) from the GE Data source are Yang, Kobayashi, and Pinkas, whereas the top three sources of data that are similar (lower $D$ values) to the GE Data source are Han, Kim, and Ejenstam.
@@
The spatial differences in the latent space can be attributed to differences in the input space, variations in the output space, or underlying fundamental differences in the input-output relationships that cannot be captured from the data. To investigate and interpret the latent spaces, a deeper study on the input and output (property) space is conducted. First, a principal component analysis (PCA) is performed to observe the differences in the input space. Figure 8c shows that the two principal components of the GE Data source cover and extend beyond the principal components of the other sources, indicating that the neither the closeness or the remoteness of the sources to the GE Data source in the latent space can be attributed to the differences in the input space. Looking further into the output distributions (Figure 8d), it is seen that the GE Data once again covers and extends beyond the properties of other available sources, demonstrating that the differences in the latent space are not due to the variabilities in the output space. As a result, the most impactful observation made from the latent space of data sources is that there are some fundamental differences between the data sources that are not explicitly known. As mentioned earlier, these possible differences for the current case can be due to differences in the manufacturing process of the alloy, testing methods, measuring procedure, underlying phenomenon, and more. Thus, it can be concluded that important underlying differences do exist between the information sources and LVGP is not only able to capture these variabilities but also learns the relationships between the sources.
@@
A further validation study is done by separating the GE Data into two different sources as GE Data 1 and GE Data 2 and retraining the LVGP model to observe the latent space again. The aim of this validation study is to observe the new latent space when the same underlying experimental procedures are implemented for two resources along with other sources that have different unknown underlying procedures. The new constructed latent space is shown in Figure 8b. A significant observation from the validation study is that the latent variables of GE Data 1 and GE Data 2 are located adjacent to each other. This result was expected since the same experimental procedures are implemented for both sources. However, the results also once again demonstrated and validated that the LVGP model is able to capture the differences in the known and unknown underlying parameters that influence the material system and learn the relationships between the sources. This remarkable capability provided by the LVGP can be highly important for further downstream tasks including but not limited to model validation, dimension reduction, sensitivity analysis, and more. It is important to highlight that without using the LVGP approach proposed in this paper, it would be difficult to learn and assess the relationships between the data sources.
@@
Figure 8: (a) The latent Space of FeCrAl sources from LVGP. GE Data is denoted as the reference source (*) for dissimilarity metric, $D$, calculations. (b) The latent Space obtained from the interpretation validation study. GE Data 1 is denoted as the reference source (*). (c) Two principal components of the input space and (d) output distribution of HV with respect to information sources
@@
4.4.4 Enhancing Modeling Capabilities Through Targeted Source Selection Based On Latent Variables